# Focus Session: LLM4PQC – An Agentic Framework for Accurate and Efficient Synthesis of PQC Cores

### v1

@@ Abstract
The design of post-quantum cryptography (PQC) hardware is a complex and hierarchical process with many challenges. A primary bottleneck is the conversion of PQC reference codes from C to high-level synthesis (HLS) specifications, which requires extensive manual refactoring [2, 19, 18]. Another bottleneck is the scalability of synthesis for complex PQC primitives, including number theoretic transform (NTT) accelerators and wide memory interfaces. While large language models (LLMs) have shown remarkable results for coding in general-purpose languages like Python, coding for hardware design is more challenging; feedback-driven and agentic integration are key principles of successful state-of-the-art approaches. Here, we propose LLM4PQC, an LLM-based agentic framework that refactors high-level PQC specifications and reference C codes into HLS-ready and synthesizable C code. Our framework generates and verifies the resulting RTL code. For correctness, we leverage a hierarchy of checks, covering fast C compilation and simulation as well as RTL simulation. Case studies on NIST PQC reference designs demonstrate a reduction in manual effort and accelerated design-space exploration compared to traditional flows. Overall, LLM4PQC provides a powerful and efficient pathway for synthesizing complex hardware accelerators.
@@ I Introduction
We present LLM4PQC, an agentic, feedback-driven LLM workflow to transform PQC reference code into synthesizable HLS-C code, considering end-to-end verification and DSE for power, performance, and/or area (PPA) objectives. Among other aspects, we focus on refactoring barriers exposed by real-world PQC kernels: HLS-incompatible constructs like calls to math.h functions, floating-point arithmetic, and runtime routines that build-up constant tables. Overall, our LLM4PQC framework integrates: (i) preprocessing transformations tailored to target HLS-C, (ii) automated C-to-HLS transformation and evaluation (using our C2HLSC [4] tool integrated with Catapult for compilation, simulation, and synthesis). Feedback from compilation/simulation failures help to iteratively refine the different transformation outcomes.
@@ IV-C HLS-C Conversion
TABLE II: Success rate, runs for C compilation and HLS, and synthesis results. HLS Synthesis is configured for latency optimization. Results are for the Nangate 45nm ASIC library.
@@
[%] Avg. Min. Max. Avg. Min. Max. Avg. Min. Max. Avg. Min. Max.
Kyber-NTT 100 14.22 10 22 7.67 6 11 2957.54 2781.7 3048.5 51.8 44 57
Dilithium-NTT 100 11.56 10 16 6.44 6 9 6506.94 6137.1 6822.2 2.8 2 3
Falcon-NTT 70 13.17 10 17 6.00 1 11 162689.64 160646.1 163163.2 15 15 15
Falcon-Sampler 50 15 12 20 7.60 6 9 113409.85 1188.87 159425.27 12240 1 20436
Falcon-FFT 100 11.67 10 18 6.67 6 10 56802.5 44702.2 62360.5 17.8 16 18
@@
Primitive Work FPGA Freq ($MHz$) $C$ LUTs $F$s DSPs BRAMs Latency ($\mu s$)
Kyber-NTT [15] Artix-7 250 277 4834 4683 0 1 1.11
LLM4PQC Artix-7 75 5639 146 119 3 1 75.00
Dilithium-NTT [15] Artix-7 180 128 7451 5275 0 0 0.71
LLM4PQC Artix-7 70 6664 370 258 5 2 95.20
Falcon-NTT* [10] Artix-7 117 4468 2119 1058 8 3 38.00
LLM4PQC Artix-7 38 30741 982 555 7 2 809.00
Falcon-FFT# [6] Artix-7 134 2048 17395 7950 20 4 15.30
LLM4PQC Artix-7 25 10768 8447 1945 33 2 430.00
Falcon-Sampler [5] XCZU7EV 300 111 7683 8266 108 0 0.73
LLM4PQC XCZU7EV 150 59 17619 3927 18 1 0.39
@@ IV-D HLS Performance: Area and Latency
Latency results in Table II show distinct characteristics across primitives. Dilithium-NTT achieves the lowest latency at 2.8 cycles on average, while Kyber-NTT requires 51.8 cycles with more consistent results (44–57 cycles). Falcon-NTT and Falcon-FFT maintain stable low latency at 15 and 17.8 cycles, respectively, across all successful conversion runs. Falcon-Sampler displays significant variation (1–20,436 cycles, with an average of 12,240), due to the data-dependent nature of rejection sampling, where lower latencies represent aggressive optimizations and higher values indicate conservative implementations preserving the complete iterative logic.
@@ IV-E RTL Simulation and Verification
Discrepancies between C-level cycle counts (Table II) and RTL results (Table III) arise because C simulation ignores pipeline overheads (fill, drain, and register delays) that are explicitly captured in RTL. Additionally, the high C-level latency for the sampler reflects the data-dependent nature of Gaussian rejection sampling, where worst-case iterations drive the metric. In contrast, RTL simulations utilize fixed deterministic seeds, capturing specific execution instances rather than worst-case behavior, resulting in lower, instance-specific cycle counts.
@@ V Discussion
Critically, our results highlight a distinct trade-off between area and latency. We observed that LLM4PQC tends to generate highly area-efficient designs, often surpassing manual baselines in LUT and FF utilization. However, this efficiency comes at the cost of higher latency. This suggests that, without explicit guidance, the LLM favors compact, sequential loop structures over the aggressive unrolling and parallelization strategies typical of hand-crafted cores. This variance confirms that coding style directly dictates hardware quality and reveals substantial opportunities for DSE at the C level. Shifting the framework’s focus from fundamental synthesizability to performance tuning, specifically targeting latency reduction through architectural refinements, remains a key frontier.
@@ VI Conclusion
We have presented LLM4PQC, an agentic framework for translation of PQC reference implementations into synthesizable hardware accelerators. By integrating LLM-driven refactoring with a compile-simulate-synthesize feedback loop, we successfully generated functional hardware for complex kernels, including the Falcon Sampler and NTT/FFT primitives. Our experimental results demonstrate that LLM4PQC effectively navigates the trade-off between design productivity and hardware quality. Our designs prove highly competitive, achieving lower resource utilization than several hand-optimized manual baselines. While this compactness currently incurs a latency penalty compared to manual designs, our approach outperforms the state-of-the-art LLM-based prior work [5] in both FF utilization and latency. This confirms that an agentic, feedback-driven workflow is superior to direct LLM translation.
Promising directions for future work include fine-tuning LLMs on HLS datasets or employing retrieval-augmented generation (RAG) to query verified HLS patterns during code generation. Additionally, we envision an LLM-driven evolutionary framework where we iteratively mutate C code (e.g., via loop restructuring or array partitioning). To accelerate this exploration, pre-synthesis PPA prediction models, similar to VeriLoC [8], could provide rapid estimates for timing and area without the latency of full synthesis runs.

### v2

# LLM4PQC - Accurate and Efficient Synthesis of PQC Cores by Feedback-Driven LLMs

@@ Abstract
The design of post-quantum cryptography (PQC) hardware is a complex and hierarchical process with many challenges. A primary bottleneck is the conversion of PQC reference codes from C to high-level synthesis (HLS) specifications, which requires extensive manual refactoring [2, 19, 18]. Another bottleneck is the scalability of synthesis for complex PQC primitives, including number theoretic transform (NTT) accelerators and wide memory interfaces. While large language models (LLMs) have shown remarkable results for coding in general-purpose languages like Python, coding for hardware design is more challenging; feedback-driven and agentic integration are key principles of successful state-of-the-art approaches. Here, we propose LLM4PQC, an LLM-based framework that refactors high-level PQC specifications and reference C codes into HLS-ready and synthesizable C code. Our framework generates and verifies the resulting RTL code. For correctness, we leverage a hierarchy of checks, covering fast C compilation and simulation as well as RTL simulation. Case studies on NIST PQC reference designs demonstrate a reduction in manual effort and accelerated design-space exploration compared to traditional flows. Overall, LLM4PQC provides a powerful and efficient pathway for synthesizing complex hardware accelerators.
@@ I Introduction
We present LLM4PQC, a feedback-driven LLM workflow to transform PQC reference code into synthesizable HLS-C code, considering end-to-end verification and DSE for power, performance, and/or area (PPA) objectives. Among other aspects, we focus on refactoring barriers exposed by real-world PQC kernels: HLS-incompatible constructs like calls to math.h functions, floating-point arithmetic, and runtime routines that build-up constant tables. Overall, our LLM4PQC framework integrates: (i) preprocessing transformations tailored to target HLS-C, (ii) automated C-to-HLS transformation and evaluation (using our C2HLSC [4] tool integrated with Catapult for compilation, simulation, and synthesis). Feedback from compilation/simulation failures help to iteratively refine the different transformation outcomes.
@@ IV-C HLS-C Conversion
TABLE II: Success rate, runs for C compilation and HLS, and synthesis results. HLS Synthesis is configured for area optimization. Results are for the Nangate 45nm ASIC library.
@@
[%] Avg. Min. Max. Avg. Min. Max. Avg. Min. Max.
Kyber-NTT 100 14.22 10 22 7.67 6 11 2,957.54 2,781.7 3,048.5
Dilithium-NTT 100 11.56 10 16 6.44 6 9 6,506.94 6,137.1 6,822.2
Falcon-NTT 70 13.17 10 17 6.00 1 11 162,689.64 160,646.1 163,163.2
Falcon-Sampler 40 15.75 12 20 7.25 6 8 141,465.1 93,961.07 159,425.27
Falcon-FFT 100 11.67 10 18 6.67 6 10 56,802.5 44,702.2 62,360.5
@@
Primitive Work FPGA Freq ($MHz$) # Cycles LUTs $F$s DSPs BRAMs Latency ($\mu s$)
Kyber-NTT [15] Artix-7 250 277 4,834 4,683 0 1 1.11
LLM4PQC Artix-7 75 5,639 146 119 3 1 75.00
Dilithium-NTT [15] Artix-7 180 128 7,451 5,275 0 0 0.71
LLM4PQC Artix-7 70 6,664 370 258 5 2 95.20
Falcon-NTT* [10] Artix-7 117 4,468 2,119 1,058 8 3 38.00
LLM4PQC Artix-7 38 30,741 982 555 7 2 809.00
Falcon-FFT# [6] Artix-7 134 2,048 17,395 7,950 20 4 15.30
LLM4PQC Artix-7 25 10,768 8,447 1,945 33 2 430.00
Falcon-Sampler [5] XCZU7EV 300 111 7,683 8,266 108 0 0.73
LLM4PQC XCZU7EV 150 59 17,619 3,927 18 1 0.39
@@ IV-D HLS Performance: Area and Latency
Latency extracted from the synthesis reports provide unreliable cycle counts, which are thus not reported. This is because several loops in the primitives’ algorithms are not statically bounded; their iteration counts cannot be correctly determined at compile time by the HLS tools. Although the actual cycle counts are not precise, their ranges and orders of magnitude still provide valuable insights. For NTT and FFT, the variation in cycle counts is small across all primitives. For the Falcon-Sampler, however, varations are significant. This is caused by the data-dependent nature of Gaussian rejection sampling, where lower latencies correspond to aggressive optimizations, while higher values represent conservative implementations that preserve the full iterative logic.
@@ V Discussion
We observed that LLM4PQC tends to generate highly area-efficient designs, often surpassing manual baselines in LUT and FF utilization. However, this efficiency comes at the cost of higher latency. This suggests that, without explicit guidance, the LLM favors compact, sequential loop structures over the aggressive unrolling and parallelization strategies typical of hand-crafted cores. This variance confirms that coding style directly dictates hardware quality and reveals substantial opportunities for DSE at the C level. Shifting the framework’s focus from fundamental synthesizability to performance tuning, specifically targeting latency reduction through architectural refinements, remains a key frontier.
@@ VI Conclusion
We have presented LLM4PQC, a novel framework for translation of PQC reference implementations into synthesizable hardware accelerators. By integrating LLM-driven refactoring with a compile-simulate-synthesize feedback loop, we successfully generated functional hardware for complex kernels, including the Falcon Sampler and NTT/FFT primitives. Our experimental results demonstrate that LLM4PQC effectively navigates the trade-off between design productivity and hardware quality. Our designs prove highly competitive, achieving lower resource utilization than several hand-optimized manual baselines. While this compactness currently incurs a latency penalty compared to manual designs, our approach outperforms the state-of-the-art LLM-based prior work [5] in both FF utilization and latency. This confirms that a feedback-driven workflow is superior to direct LLM translation.
Promising directions for future work include closer/agentic integration of the LLM-driven phases in LLM4PQC, fine-tuning the LLMs on HLS datasets, or employing retrieval-augmented generation (RAG) to query verified HLS patterns during code generation. Additionally, we envision an LLM-driven evolutionary framework where we iteratively mutate and explore C codes for ease of hardware generation (e.g., via loop restructuring or array partitioning). To accelerate such exploration, pre-synthesis PPA prediction models, similar to VeriLoC [8], could provide rapid estimates for timing and area without the latency of full synthesis runs.
VII Acknowledgments
Buddhi Perera and Ramesh Karri were supported in part by the National Science Foundation under Grant No. #2450539. The authors acknowledge support from the NYU Center for Cybersecurity (CCS) and the NYU Abu Dhabi Center for Cybersecurity (CCS-NYUAD).